The first three terms of a geometric progression are 100, 90 and 81.
Find the common ratio of this progression.?

Found 2 solutions by Boreal, stanbon:
Answer by Boreal(15235)

The common ratio: 90/100=0.9
Try it for the next one: 81/90=0.9
The common ratio is 0.9. The next term will be 72.9

Answer by stanbon(75887)

first three terms of a geometric progression are 100, 90 and 81.
Find the common ratio of this progression.
-----
r = 90/100 = 0.9
---------------------
Cheers,
Stan H.
----